HELP!!! Depreciation Method: straight line Purchase Date: 8/1/2005 Cost: $20,000.00 Estimated Life: 5 Sales Price: $1,000.00...

HELP!!!

Depreciation Method: straight line

Purchase Date: 8/1/2005

Cost: $20,000.00

Estimated Life: 5

Sales Price: $1,000.00

Date Sold: 12/31/2010

a. What is the depreciation expense for yr 3 using the straight‐line method?

b. What is the net book value of the equipment at the end of year 2?

c. Calculate gain/loss on sale of equipment.

d. What is the account balance in accumulated depreciation on 12/31/2010?
Answer #1

a.

Annual depreciation expense = Cost of equipment / Useful life

= 20,000/5

= $4,000

Hence, Depreciation expense for year 3 = $4,000

b.

Depreciation expense for 2005 = 4,000 x 5/12

= $1,667

Accumulated depreciation at the end of year 2 = 1,667+4,000

= $5,667

Book value of equipment at the end of year 2 = 20,000-5,667

= $14,333

d.

Accumulated depreciation on 12/31/2010 = $1,667 for 2005 + $4,000 for 2006 + $4,000 for 2007 + $4,000 for 2008 + $4,000 for 2009 + $2,333 for 2010

= $20,000

c.

Book value of equipment on 12/31/2010 = Cost price - Accumulated depreciation on 12/31/2010

= 20,000-20,000

= $0

Gain on sale of equipment = Sale price - Book value of equipment on 12/31/2010

= 1,000-0

= $1,000

Kindly comment if you need further assistance. Thanks‼!